# Mixing Artificial and Natural Intelligence: From Statistical Mechanics to AI and Back to Turbulence

### v1

@@ Abstract
The paper reflects on the future role of AI in scientific research, with a special focus on turbulence studies, and examines the evolution of AI, particularly through Diffusion Models rooted in non-equilibrium statistical mechanics. It underscores the significant impact of AI on advancing reduced, Lagrangian models of turbulence through innovative use of deep neural networks. Additionally, the paper reviews various other AI applications in turbulence research and outlines potential challenges and opportunities in the concurrent advancement of AI and statistical hydrodynamics. This discussion sets the stage for a future where AI and turbulence research are intricately intertwined, leading to more profound insights and advancements in both fields.
@@
I begin in Section 1 by defining AI, highlighting key components such as Automatic Differentiation, Deep Learning, Reinforcement Learning, and Generative Models, and discussing their progressive interdependence. This preliminary discussion also helps to set the stage for methodologies of how we use AI (as System 1 or System 2) and explain how AI’s capacity to navigate the curse of dimensionality has sparked significant interest in scientific communities, enabling novel approaches to longstanding computational challenges.
@@
Section 4.6 is devoted to our recent exploration of the diffusion models of AI – a class of generative AI models that is based on non-equilibrium statistical mechanics, which have remarkable abilities to generate high-fidelity simulations from learned data distributions, and shares many common features, which are far from being explored, with turbulence.
In Section 5, we pivot towards turbulence/physics, initiating with a primer on Lagrangian Turbulence, where we trace the trajectories of fluid particles. We advocate for the Lagrangian method as a bridge linking AI’s diffusion models with turbulence studies, providing deeper insights into the intricate, multi-scale dynamics of turbulence. This approach, beneficial in pre-AI turbulence research, led to the development of a Lagrangian closure model encapsulated by a tetrad that represents the dynamics of a fluid element.
@@ 1.3 The AI Landscape
  • Automatic Differentiation: This involves speeding up the basic elements of analog computation using a set of elementary operations and the chain rule. It enables accurate and efficient gradient evaluations for optimization and machine learning. (For a comprehensive review, consult Baydin et al. (2015).)
@@ 2.3 Engineering Sciences
On the application side of engineering sciences, I am predominantly interested in the study of energy system networks, including power Bienstock et al. (2014); Turitsyn et al. (2011); Dörfler et al. (2013), natural gas Chertkov et al. (2014); Zlotnik et al. (2016), and district heating-cooling systems Métivier & Chertkov (2020); Chertkov & Novitsky (2019); Valenzuela et al. (2023). While working on these topics I am focusing on understanding, proposing novel approaches, and developing algorithms, including AI-based  [^1] Here are highlights of my contributions in Power System Informed AI, thus far predominantly via Deep Neural Networks: Deka et al. (2017); Lokhov et al. (2018b); Stulov et al. (2020); Pagnier & Chertkov (2021b; a); Afonin & Chertkov (2021); Pagnier et al. (2022); Ferrando et al. (2024). A review of these efforts is forthcoming in a separate article I intend to write., for constructing, reinforcing, managing, and controlling these energy systems. This involves accounting for customer and operator behavioral factors and integrating the energy system seamlessly with other critical infrastructures, including transportation.
@@ 2.4 The Science of AI
    Diffusion Models of AI: Quite a lot of my recent focus is on diffusion models – a remarkable tool in generative AI. These models are renowned for their ability to generate synthetic images based on textual prompts or stated in a bit more streamlined form – training set of images, considered as i.i.d. samples from a probability distribution. What makes the diffusion models even more intriguing is their foundation in the clever utilization of stochastic ordinary differential equations (ODEs) and time-inversion principles, stemming from non-equilibrium statistical mechanics and the aforementioned stochastic optimal control. We will be discussing more of this in Sections 4.4,4.6,4.7,4.8.
@@
    AI Transformers: AI Transformers: As a standout generative AI tool, transformers form the foundation of recent groundbreaking advancements, including sophisticated Large Language Models such as ChatGPT. Transformers operate incrementally, extending their output word-by-word (or token-by-token, particle-by-particle) from a given input or query. Yet, at each step, the ensemble of existing particles – comprising both the output generated and the input provided thus far – undergoes a dynamic evolution, effectively converging on a lower-dimensional manifold or distribution that informs the sampling of the subsequent particle. As suggested by Geshkovski et al. (2023), the impressive efficacy of transformer models can be ascribed to their function as dynamical systems, particularly how they leverage the tendency of multi-particle systems to demonstrate clustering or exhibit chaotic dynamics. My keen interest lies in exploring whether the multi-particle (Lagrangian) interpretation of turbulent flows discussed in Section 5.5 could enlighten the design of more refined transformers and, by extension, generative models at large.
@@ 3.1 Physics-Blind AI discovers Physics
In the turbulence research community, AI tools have been increasingly applied, specifically Deep Neural Networks within supervised learning (treating turbulence as an input-output function) or generative models for handling spatial or spatio-temporal snapshots of turbulence. These AI tools have been employed to process these snapshots, treating them as if they were images of celebrities or segments from movies. This approach can be likened to a System 1 utilization of AI in turbulence research.
@@ 3.2 Exploring Physics-Informed AI
    Understanding via Reduced Modeling: Consider the aforementioned curse of dimensionality, a concern in many physical disciplines heavily reliant on computations. Even when we have an equation believed to be universal and accurate for various phenomena, such as the Navier-Stokes equation in fluid mechanics, running it in challenging cases, like at high Reynolds numbers, becomes prohibitively expensive. This is where model reduction comes into play. We postulate reduced models with significantly fewer degrees of freedom (for example changing from PDE to ODE), but this introduces a multitude of options for specifying these reduced models, laden with uncertainty. Here’s where physical hypotheses come handy. We discuss hypotheses that can be parameterized in terms of physically meaningful parameters (as mentioned earlier). Alternatively, we can use Neural Networks and other AI tools to express degrees of freedom about which we have limited knowledge.
Importantly, all of the above relies on data we can absolutely trust – the so-called Ground Truth (GT) data. Traditional high-resolution numerical simulations which are abundant in physics-intensive disciplines, such as Direct Numerical Simulations (DNS) commonly used in turbulence research, can serve as this GT data.
@@
Furthermore, the method initially described as "Artificial neural networks for solving ordinary and partial differential equations" Lagaris et al. (1998) has undergone a resurgence as Physics-Informed Neural Networks (PINNs) Raissi et al. (2017a; b); Karniadakis et al. (2021)  [^2] It was also termed in Karniadakis et al. (2021), somehow more pretentiously, as the Physics-Informed Machine Learning (PIML). In my (the author) opinion, and given the discussion of the pre-history above, these terms – PINN and PIML, and especially the latter one used in the literature extensively prior to the aforementioned papers, in particular to name a series of Los Alamos NL workshops – are much more general and including many other, e.g. equation-free and phenomenology guided, physics informed approaches.. PINNs operate on the premise that the underlying ODEs or PDEs are precisely known, subsequently enhancing computational techniques – historically, the collocation-point method – by integrating the equations into the loss function for DNN training. While PINN models often yield faster results than direct numerical solutions, they may do so at the cost of some accuracy.
@@ 3.3 AI to Discover New Laws of Physics
In this context, it’s crucial to embrace the spirit of bold hypothesis formation. Gone are the days of discarding seemingly unconventional ideas. With the advent of AI techniques, particularly our newfound ability to optimize in high dimensions (thanks to automatic differentiation), we can adopt a Bayesian approach and consider multiple hypotheses concurrently, possibly guided by some intuitive priors.
This shift allows us to move away from subjective hypothesis selection and rely more on data-driven approaches. AI’s capacity to assess and choose more scientifically reliable hypotheses based on data offers invaluable guidance. It’s important to note that this approach may not provide proof but serves as an additional tool to steer our attention towards what’s worth investigating and what may be disregarded.
@@ 4.1 Markov Chain Monte Carlo
where $\bm{w}(t)$ represents the Wiener process with unit covariance matrix in $d$ dimensions, and $\bm{\psi}(\bm{x})\in\mathbb{R}^{d}$ denotes what we will call (following AI jargon) the score function. Equilibrium statistical mechanics informs us that the MCMC, as defined by Eq. (1), will ultimately converge in a statistical sense. This implies that, given a set of outputs $\bm{x}(t_{1}),\ldots,\bm{x}(t_{S})$ at successive times $t_{1},\ldots,t_{S}$ – each separated by at least the so-called mixing time – the sampled data points will be i.i.d.
@@ 4.2 Annealing
The technique of annealing has been shown to be effective, offering not just empirical success but also providing a theoretical basis for constructing proofs for sampling complex distributions, as demonstrated in Jerrum & Sinclair (1989). However, annealing is also known for its slow convergence; the time required to produce subsequent i.i.d. samples is not only significant but also scales unfavorably with the size of the system.
@@ 4.5 Generating Synthetic Samples
Figure 1: GPT-4 prompt: Illustration of the University of Arizona’s bobcat mascot swimming butterfly style in a turbulent river.
@@ 4.6 AI Diffusion Algorithm
which yields the marginal probability as a Gaussian/normal mixture over the Ground Truth (GT) samples:
@@
Herein lies the transformative power of AI: we approximate the score function using a deep neural network (NN) parameterized by ${\bm{\theta}}$. The training process involves optimizing the discrepancy between the true score function and its NN approximation, evaluated over a uniform temporal distribution within $[0,T]$ and spatial distribution according to Eq. (6):
@@
With this training completed, we can now generate new samples from the distribution represented by its GT samples. We simply initialize the reverse process (7) with a draw from the univariate normal distribution ${\bm{x}}(T)\sim\mathcal{N}(0,{\bm{I}})$ and consider ${\bm{x}}(0)$ as the new synthetic sample. Remarkably, although these new samples are distinct, they bear a "similarity" to the GT samples – a testament to the deep neural network’s capacity to effectively smooth the distribution around $t=0$ (where the empirical distribution turns simply into a sum of the $\delta$-functions).
@@ 4.7.1 The U-Turn Technique
Figure 2: Synthetic images generated from the model presented in Behjoo & Chertkov (2023), starting at $t=0$, and executing U-turns at various times within the $[0,1000]$ interval. The model was trained on a set of 1000 ground truth (GT) butterfly images sourced from the Hugging Face Hub (https://huggingface.co/datasets/huggan/smithsonian_butterflies_subset). Early U-turns, close to $t=0$, yield images resembling the original GT images. Conversely, U-turns performed later, specifically at $t\gtrsim 600$, result in the creation of novel synthetic images.
The U-turn approach, as introduced in Behjoo & Chertkov (2023), explores the pivotal insight that the core measure of the diffusion models’ efficiency lies in their ability to rapidly attenuate correlations during the reverse or de-noising phase. Contrary to extending the forward diffusion to its full extent, the U-turn technique advocates for an earlier reversal - illustrated in Fig. (2). Specifically, it employs a shortened forward diffusion phase, immediately followed by the customary reverse dynamics, which commence from the terminal state of the forward phase. To calibrate the optimal timing for the U-turn, the methodology employs a suite of analytical tools, including auto-correlation analysis, the weighted norm of the score function, and the Kolmogorov-Smirnov test for Gaussianity.
@@ 4.7.3 Space-Time Diffusion
4.7.4 Phase Transitions in Optimally Trained AI Diffusion Models
In the recent work Biroli et al. (2024) a detailed analysis inspired by the concept of phase transitions in physics is applied to optimally trained AI diffusion models in the limits of infinite dimensions and sample sizes $d,S\to\infty$. This study commences with an analytical exploration based on large-deviation (saddle-point) analysis, focusing on sampling from a synthetic distribution characterized by a sum of two Gaussian distributions. Subsequent experimental validations on a pre-trained AI diffusion model for realistic high-dimensional data set of images corroborated the theoretical insights, unveiling three distinct dynamical regimes that emerge along the reverse generative path of the diffusion process.

### v2

@@ Abstract
The paper reflects on the future role of AI in scientific research, with a special focus on turbulence studies, and examines the evolution of AI, particularly through Diffusion Models rooted in non-equilibrium statistical mechanics. It underscores the significant impact of AI on advancing reduced, Lagrangian models of turbulence through innovative use of DNNs. Additionally, the paper reviews various other AI applications in turbulence research and outlines potential challenges and opportunities in the concurrent advancement of AI and statistical hydrodynamics. This discussion sets the stage for a future where AI and turbulence research are intricately intertwined, leading to more profound insights and advancements in both fields.
@@
I begin in Section 1 by defining AI, highlighting key components such as Automatic Differentiation, Deep Learning, Reinforcement Learning, and Generative Models, and discussing their progressive interdependence 111 In this manuscript, AI is used as a collective term to encompass various techniques, including but not limited to Machine Learning (ML). ML is a subset of AI focused on algorithms that enable computers to learn from and make predictions based on data. This preliminary discussion also helps to set the stage for methodologies of how we use AI (as System 1 or System 2) and explain how AI’s capacity to navigate the curse of dimensionality has sparked significant interest in scientific communities, enabling novel approaches to longstanding computational challenges.
@@
Section 4.6 is devoted to our recent exploration of the diffusion models – a class of generative AI models that is based on non-equilibrium statistical mechanics, which have remarkable abilities to generate high-fidelity simulations from learned data distributions, and shares many common features, which are far from being explored, with turbulence.
In Section 5, we pivot towards turbulence/physics, initiating with a primer on Lagrangian Turbulence, where we trace the trajectories of fluid particles. We advocate for the Lagrangian method as a bridge linking diffusion models with turbulence studies, providing deeper insights into the intricate, multi-scale dynamics of turbulence. This approach, beneficial in pre-AI turbulence research, led to the development of a Lagrangian closure model encapsulated by a tetrad that represents the dynamics of a fluid element.
@@ 1.3 The AI Landscape
  • Automatic Differentiation: often abbreviated as AD, is a set of techniques to evaluate the derivative of a function specified by a computer program. Unlike numerical differentiation (which approximates derivatives using finite differences and can be prone to numerical errors) or symbolic differentiation (which manipulates complex mathematical expressions to find derivatives and can be computationally intensive), AD provides a way to calculate derivatives exactly and efficiently by systematically applying the chain rule of calculus. (For a comprehensive review, consult Baydin et al. (2015); Lavin et al. (2022).)
@@ 2.3 Engineering Sciences
On the application side of engineering sciences, I am predominantly interested in the study of energy system networks, including power Bienstock et al. (2014); Turitsyn et al. (2011); Dörfler et al. (2013), natural gas Chertkov et al. (2014); Zlotnik et al. (2016), and district heating-cooling systems Métivier & Chertkov (2020); Chertkov & Novitsky (2019); Valenzuela et al. (2023). While working on these topics I am focusing on understanding, proposing novel approaches, and developing algorithms, including AI-based  [^2] Here are highlights of my contributions in Power System Informed AI, thus far predominantly via DNNs: Deka et al. (2017); Lokhov et al. (2018b); Stulov et al. (2020); Pagnier & Chertkov (2021b; a); Afonin & Chertkov (2021); Pagnier et al. (2022); Ferrando et al. (2024). A review of these efforts is forthcoming in a separate article I intend to write., for constructing, reinforcing, managing, and controlling these energy systems. This involves accounting for customer and operator behavioral factors and integrating the energy system seamlessly with other critical infrastructures, including transportation.
@@ 2.4 The Science of AI
    Diffusion Models: Quite a lot of my recent focus is on diffusion models – a remarkable tool in generative AI. These models are renowned for their ability to generate synthetic images based on textual prompts or stated in a bit more streamlined form – training set of images, considered as i.i.d. samples from a probability distribution. What makes the diffusion models even more intriguing is their foundation in the clever utilization of stochastic ordinary differential equations (ODEs) and time-inversion principles, stemming from non-equilibrium statistical mechanics and the aforementioned stochastic optimal control. We will be discussing more of this in Sections 4.4,4.6,4.7,4.8.
@@
    Transformers: As a standout generative AI tool, transformers introduced in Vaswani et al. (2017) form the foundation of recent groundbreaking advancements, including sophisticated Large Language Models such as ChatGPT. Transformers operate incrementally, extending their output word-by-word (or token-by-token, particle-by-particle) from a given input or query. Yet, at each step, the ensemble of existing particles – comprising both the output generated and the input provided thus far – undergoes a dynamic evolution, effectively converging on a lower-dimensional manifold or distribution that informs the sampling of the subsequent particle. As suggested by Geshkovski et al. (2023), the impressive efficacy of transformer models can be ascribed to their function as dynamical systems, particularly how they leverage the tendency of multi-particle systems to demonstrate clustering or exhibit chaotic dynamics. My keen interest lies in exploring whether the multi-particle (Lagrangian) interpretation of turbulent flows discussed in Section 5.5 could enlighten the design of more refined transformers and, by extension, generative models at large.
@@ 3.1 Physics-Blind AI discovers Physics
In the turbulence research community, AI tools have been increasingly applied, specifically DNNs within supervised learning (treating turbulence as an input-output function) or generative models for handling spatial or spatio-temporal snapshots of turbulence. These AI tools have been employed to process these snapshots, treating them as if they were images of celebrities or segments from movies. This approach can be likened to a System 1 utilization of AI in turbulence research.
@@ 3.2 Exploring Physics-Informed AI
    Understanding via Reduced Modeling: Consider the aforementioned curse of dimensionality, a concern in many physical disciplines heavily reliant on computations. Even when we have an equation believed to be universal and accurate for various phenomena, such as the Navier-Stokes equation in fluid mechanics, running it in challenging cases, like at high Reynolds numbers, becomes prohibitively expensive. This is where model reduction comes into play. We postulate reduced models with significantly fewer degrees of freedom (for example changing from Partial Differential Equations (PDE) to Ordinary Differential Equations (ODE)), but this introduces a multitude of options for specifying these reduced models, laden with uncertainty. Here’s where physical hypotheses become handy. We discuss hypotheses that can be parameterized in terms of physically meaningful parameters (as mentioned earlier). Alternatively, we can use Neural Networks and other AI tools to express degrees of freedom about which we have limited knowledge.
Importantly, all of the above relies on data we can absolutely trust – the so-called Ground Truth (GT) data.
The term GT in this manuscript refers to the high-fidelity data used for training and validating AI models. In the context of turbulence simulations discussed in Section 5, GT typically consists of high-fidelity Direct Numerical Simulation (DNS) data. For diffusion models, discussed in Section 4, GT refers to the ensemble of real (not synthetic) images that the model aims to replicate. It is important to note that while GT data is considered a standard reference, it may contain errors due to numerical approximations or human factors. These errors can affect the training and performance of AI models. Acknowledging and modeling these errors is crucial for improving the robustness of AI methods. Traditional approaches in AI and ML, such as Graphical Models, have been employed to account for errors and uncertainties in the data (notably to de-noise images, see e.g., Koller & Friedman (2009)), and similar methods can be extended to modern AI techniques.
@@
Furthermore, the method initially described as "Artificial neural networks for solving ordinary and partial differential equations" Lagaris et al. (1998) has undergone a resurgence as Physics-Informed Neural Networks (PINNs) Raissi et al. (2017a; b); Karniadakis et al. (2021). It was also termed in Karniadakis et al. (2021), somehow more pretentiously, as Physics-Informed Machine Learning (PIML). In the author’s opinion, and given the discussion of the pre-history above, these terms – PINN and PIML, and especially the latter one used in the literature extensively prior to the aforementioned papers, in particular to name a series of Los Alamos NL workshops on Physics Informed Machine Learning conducted in 2016, 2018, 2020, and 2022 in Santa Fe and with the next one scheduled for October of 2024 – are much more general and inclusive, encompassing many other approaches, such as equation-free and phenomenology-guided physics-informed methods. PINNs operate on the premise that the underlying ODEs or PDEs are precisely known, subsequently enhancing computational techniques – historically, the collocation-point method – by integrating the equations into the loss function for DNN training. While PINN models often yield faster results than DNS, they may do so at the cost of some accuracy.
@@ 3.3 AI to Discover New Laws of Physics
In this context, it’s crucial to embrace the spirit of bold hypothesis formation. Gone are the days of discarding seemingly unconventional ideas. With the advent of AI techniques, particularly our newfound ability to optimize in high dimensions (thanks to AD), we can adopt a Bayesian approach and consider multiple hypotheses concurrently, possibly guided by some intuitive priors.
This shift allows us to move away from subjective hypothesis selection and rely more on data-driven approaches. AI’s capacity to assess and choose more scientifically reliable hypotheses based on data offers invaluable guidance. It’s important to note that this approach may not provide proof but serves as an additional tool to steer our attention towards what’s worth investigating and what may be disregarded. We briefly explore some intriguing steps in this direction in Section 5.6.
@@ 4.1 Markov Chain Monte Carlo
where $\bm{w}(t)$ represents the Wiener process with unit covariance matrix in $d$ dimensions, and $\bm{\psi}(\bm{x})\in\mathbb{R}^{d}$ denotes what we will call (following originally stiatistics, and then AI jargon) the score function 333 The term ”score function” originates from statistics, dating back to the foundational work of Ronald Fisher; see e.g., Wikipedia . It has been widely used in computational statistics, particularly in methods such as MCMC and Hamiltonian Monte Carlo. In the context of machine learning, it has been employed in Hyvärinen (2005). More recently, the role of the score function in diffusion models was clearly described in Song & Ermon (2020). We thank the anonymous reviewer for highlighting these important references and providing clarity on the historical context of the term. . Equilibrium statistical mechanics informs us that the MCMC, as defined by Eq. (1), will ultimately converge in a statistical sense. This implies that, given a set of outputs $\bm{x}(t_{1}),\ldots,\bm{x}(t_{S})$ at successive times $t_{1},\ldots,t_{S}$ – each separated by at least the so-called mixing time – the sampled data points will be i.i.d.
@@ 4.2 Annealing
The technique of annealing has been shown to be effective, offering not just empirical success but also providing a theoretical basis for constructing proofs for sampling complex distributions, as demonstrated in Jerrum & Sinclair (1989). However, annealing is also known for its slow convergence; the time required to produce subsequent i.i.d. samples is not only significant but also scales unfavorably with the size of the system  [^4] The issue of the big-O complexity of the annealing scheme is resolved mathematically through the so-called Fully Polynomial Randomized Approximation Schemes (FPRAS). These schemes are developed for a very special class of relatively simple models, such as attractive (ferromagnetic) Ising models. The complexity of even the best FPRASs is polynomial. For example, for certain instances of the Ising model, the time complexity can be $O(n^{2}\log\frac{1}{\epsilon})$, where $n$ is the number of nodes in the graph and $\epsilon$ is the approximation parameter. See Jerrum & Sinclair (1989) for additional details..
@@ 4.5 Generating Synthetic Samples
Figure 1: GPT-4 prompt: Illustration of the University of Arizona’s bobcat mascot swimming butterfly style in a turbulent river. (The entire prompt was used to generate the image without providing additional information such as color, symmetry, or viewing direction.)
@@ 4.6 Diffusion Algorithm
which yields the marginal probability as a Gaussian/normal mixture over the GT samples:
@@
Herein lies the transformative power of AI: we approximate the score function using a DNN parameterized by ${\bm{\theta}}$. The training process involves optimizing the discrepancy between the true score function and its NN approximation, evaluated over a uniform temporal distribution within $[0,T]$ and spatial distribution according to Eq. (6):
@@
With this training completed, we can now generate new samples from the distribution represented by its GT samples. We simply initialize the reverse process (7) with a draw from the univariate normal distribution ${\bm{x}}(T)\sim\mathcal{N}(0,{\bm{I}})$ and consider ${\bm{x}}(0)$ as the new synthetic sample. Remarkably, although these new samples are distinct, they bear a "similarity" to the GT samples – a testament to the DNN’s capacity to effectively smooth the distribution around $t=0$ (where the empirical distribution turns simply into a sum of the $\delta$-functions).
@@ 4.7.1 The U-Turn Technique
To illustrate the forward and reverse processes, and their use in training and generating new samples, we provide an example introduced in Behjoo & Chertkov (2023) and called the U-turn diffusion model. This model explores diffusion processes consisting of a forward noise-injecting stage and a reverse de-noising stage to encode information about the GT samples.
The U-turn diffusion model modifies the traditional approach by shortening both the forward and reverse processes, starting the reverse dynamics from the final configuration of the forward process. This modification aims to optimize the generation of synthetic samples that are independent and identically distributed (i.i.d.) according to the probability distribution represented by the GT samples.
In the experiments of Behjoo & Chertkov (2023) with the ImageNet-64 dataset Deng et al. (2009), the U-turn diffusion model achieved state-of-the-art Fréchet Inception Distance scores with fewer Neural Function Evaluations. Notably, this approach provided a 1.35-fold speed-up in inference without requiring retraining.
Figure 2 illustrates the basic Score Based Diffusion process and the U-turn concept. The analysis reported in Behjoo & Chertkov (2023), based on new tests such as the U-Turn Auto-Correlation function test, revealed that making the U-turn earlier is beneficial but should not be done too early to avoid memorization of the original image. These tests guide the optimal timing for the U-turn, ensuring the generation of high-quality synthetic images.
Figure 2: Illustration, from Behjoo & Chertkov (2023), of the basic Score Based Diffusion (left) and U-turn (right) concepts. The U-turn involves an earlier transition from the forward to the reverse process, optimizing sample generation.
@@ 4.7.3 Space-Time Diffusion
4.7.4 Phase Transitions in Optimally Trained Diffusion Models
In the recent work Biroli et al. (2024) a detailed analysis inspired by the concept of phase transitions in physics is applied to optimally trained diffusion models in the limits of infinite dimensions and sample sizes $d,S\to\infty$. This study commences with an analytical exploration based on large-deviation (saddle-point) analysis, focusing on sampling from a synthetic distribution characterized by a sum of two Gaussian distributions. Subsequent experimental validations on a pre-trained diffusion model for realistic high-dimensional data set of images corroborated the theoretical insights, unveiling three distinct dynamical regimes that emerge along the reverse generative path of the diffusion process.